# Agent Journey Beyond RGB: Unveiling Hybrid Semantic-Spatial Environmental Representations for Vision-and-Language Navigation

### v4

@@ 3.4 Hybrid Representation Aggregation
Fusing Hybrid Semantic-Spatial Environmental Representations. As depicted in Fig. 4, we employ attention pooling to project all environmental representations at time step $t$, along with the corresponding instruction, into an equivalent dimension $d$. Specifically, the attention pooling AP operation is defined as follows:
^$\hat{\mathcal{H}}=\text{AP}(\mathcal{H},v),\quad\mathcal{H}\in\{\tilde{\mathcal{X}^{\mathcal{S}}},\tilde{\mathcal{T}^{\mathcal{D}}},\tilde{\mathcal{X}^{R}},\tilde{\mathcal{T}^{R}},\mathcal{I}\},$ (4)
where $v\in\mathbb{R}^{d}$ denoting the learnable query vector and ^$\hat{\mathcal{H}}\in\mathbb{R}^{d}$. The AP operator is given by:
@@
$\small\footnotesize\mathcal{B}=\begin{bmatrix}\beta_{1}&\beta_{2}&\beta_{3}&\beta_{4}\end{bmatrix}^{\top}=\sigma(\mathrm{FFN}[\tilde{\mathcal{X}_{0}^{\mathcal{S}}};\tilde{\mathcal{X}_{0}^{\mathcal{R}}};\tilde{\mathcal{T}_{0}^{\mathcal{D}}};\tilde{\mathcal{T}_{0}^{\mathcal{R}}}])$ (6)
$\small\footnotesize E^{\mathrm{hyb}}=\begin{bmatrix}\hat{\mathcal{X}^{\mathcal{S}}}&\hat{\mathcal{X}^{\mathcal{R}}}&\hat{\mathcal{T}^{\mathcal{D}}}&\hat{\mathcal{T}^{\mathcal{R}}}\end{bmatrix}\cdot\mathcal{B},$ (7)
where $\sigma$ stands for the sigmoid activation function, and [;][;][ ; ] denotes concatenation.
@@
We convert local predictions into the global action space and predict action $a_{t}$ via dynamic fusing [6] $P_{\mathcal{T}}$ and $P_{\mathcal{X}}$.
3.5 Training and Inference
Pretraining. Our pre-training phase followed prior work [6, 30], utilizing auxiliary tasks such as Masked Language Modeling (MLM), Masked Region Classification (MRC), and Single-step Action Prediction (SAP). Object Grounding (OG) task to localize target objects in REVERIE and SOON. Given the sparsity of depth images and textual semantics compared to RGB images, completely pre-train all parameter of SUSA may lead to overfitting and increased training costs. As shown in Fig. 2, only the computation flows indicated by black arrows were pre-trained. The effectiveness of this partial pre-trained strategy is demonstrated in Sec. 4.3.
Methods Validation Seen Validation Unseen Test Unseen
[…]
RCM [54] 10.70 23.33 21.82 16.23 15.36 11.98 9.29 6.97 4.89 3.89 10.60 7.84 6.67 3.67 3.14
HOP+ [43] 10.59 55.87 49.55 40.76 36.22 14.57 36.07 31.13 22.49 19.33 15.17 33.82 28.24 20.20 16.86
DSRG [50] - 75.69 68.09 61.07 54.72 - 47.83 34.02 32.69 23.37 - 54.04 37.09 32.49 22.18
GridMM [56] - - - - - 23.20 51.37 36.47 34.57 24.56 19.97 53.13 36.60 34.87 23.45
AZHP [58] 13.60 70.98 62.24 56.99 50.14 22.08 49.02 36.25 32.41 24.13 21.10 52.52 36.11 32.10 22.54
FDA [18] - - - - - 19.04 47.57 35.90 32.06 24.31 17.30 49.62 36.45 30.34 22.08
CONSOLE [34] - 74.14 65.15 60.08 52.69 - 50.07 34.40 34.05 23.33 - 55.13 37.13 33.18 22.25
KERM [31] 14.25 71.89 64.04 57.55 51.22 21.85 49.02 34.83 33.97 24.14 18.38 52.26 37.46 32.69 23.15
VER [36] 16.13 75.83 66.19 61.71 56.20 23.03 55.98 39.66 33.71 23.70 24.74 56.82 38.76 33.88 23.19
DUET [6] 13.86 71.75 63.94 57.41 51.14 22.11 46.98 33.73 32.15 23.03 21.30 52.51 36.06 31.88 22.06
SUSA (Ours) 14.60 76.95 69.07 61.77 55.86 22.59 51.75 38.86 35.02 26.56 17.86 54.39 41.54 36.11 27.31
Table 1: Comparison with the state of the art on REVERIE. Bold highlight the best performance in each column, $\uparrow$ indicates better performance with higher values.‘–’: unavailable statistics.
Fine-tuning and Inference. During fine-tuning, we employ a contrastive learning loss, $\mathcal{L}_{cl}$ to align the hybrid environmental representations $E^{hyb}$ with the corresponding instructions $\mathcal{I}$:
@@ 3.5 Training and Inference
where $B$ and $t$ represent the batch size and the temperature. Then, we train the model based on hybrid environment representations by performing single-step action prediction [5, 6] with a contrastive learning loss $\mathcal{L}_{cl}$. The hybrid single action prediction loss in behavior cloning is computed as $\mathcal{L}_{hsap}^{gt}=\sum_{t=1}^{T}-\log p(a_{t}^{gt}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$ and $\mathcal{L}_{hsap}^{\star}=\sum_{t=1}^{T}-\log p(a_{t}^{\star}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$, where $a_{t}^{gt}$ represents the ground truth action, and $a_{t}^{\star}$ denotes the pseudo-target action sampled from current exploration map. The final training loss is then expressed as follows:
@@
During the inference phase, the agent is tasked with sequentially navigating to the predicted navigable nodes, following the shortest path algorithm [13], without making any abrupt “jumps” between nodes.
@@ 4.1 Task Setup and Implementation
Datasets. We evaluate our model on three diverse VLN benchmarks based on the Matterport3D simulator environment [2]. R2R [2] focuses solely on navigation following detailed instructions. REVERIE [41] requires the agent to recognize the correct object from candidate bounding boxes upon reaching the navigation goal. SOON [64] further challenges agents to generate candidate bounding boxes using an object detector, as predefined boxes are not provided.
Metrics. We evaluate the navigation performance using standard metrics: Trajectory Length (TL), Success Rate (SR), Navigation Error (NE), and Success Rate weighted by Path Length (SPL). Additionally, we use Normalized Dynamic Time Warping (nDTW) and Success weighted by Normalized Dynamic Time Warping (sDTW) to evaluate the fidelity of successful trajectories to their corresponding instructions. For the REVERIE and SOON tasks, which involve object identification, we also evaluate object grounding metrics: Remote Grounding Success (RGS), and RGS weighted by Path Length (RGSPL). Given that SPL optimally balances navigation success rate and trajectory length, we employ it as our primary metric.
Implementation Details. We initialized transformer layers with pretrained LXMERT [48] during pre-training. Pre-training was performed for 100k iterations with a batch size of 32. Fine-tuning was performed for 25k iterations. For fine-tuning, we used batch sizes of 4, 8, and 2 for R2R, REVERIE, and SOON, respectively. The default hyperparameters are $\lambda_{1}$ = 0.2 and $\lambda_{2}$= 0.8. The node and view dimensions are identical: $d_{v}=d_{n}=d=768$. We minimized changes to the baseline (DUET) [6] settings and used no additional annotations. All experiments were conducted on a single NVIDIA RTX 4090 GPU. Detailed experimental setup is provided in the supplementary material.
@@ 4.2 Comparisons with State of the Art
We conduct a comprehensive comparison on the REVERIE [41], R2R [2], and SOON [64] datasets. For fairness, we exclude methods that involve pre-exploration [66, 46] or utilize large-scale data augmentation [7, 55]. REVERIE. Table 1 presents a comparative analysis on the REVERIE. On the test unseen split, SUSA achieves an SPL score of 41.54%, significantly outperforming the previous state-of-the-art (SOTA) method by a large margin of 2.78%. Furthermore, we establish new SOTA results for RGS and RGSPL on both the validation and test unseen splits, demonstrating that the introduced textual semantics substantially elevate overall navigation, particularly benefiting object localization performance. R2R. The compared experiment results on R2R, summarized in  Tab. 2. Compared with the SOTA ESCENE [62] model on the validation and test unseen splits, SUSA achieves 64.85% and 63.83% on SPL, respectively. These gains manifest not only an improved navigation performance but also enhanced navigation efficiency, since SPL considers both success rate and trajectory length.
Methods Validation Unseen Test Unseen
[…]
RCM [54] 11.46 43 - 6.09 11.97 43 38 6.12
HOP [42] 12.27 64 57 3.80 12.68 64 59 3.83
DSRG [50] - 73 62 3.00 - 72 61 3.33
PanoGen [28] 13.40 74 64 3.03 14.38 71 61 3.31
GridMM [56] - 75 64 - - 73 62 -
FDA [18] 13.68 72 64 3.41 14.76 69 62 3.41
KERM [31] 13.54 71.95 60.91 3.22 14.60 69.73 59.25 3.61
AZHP [58] 13.68 71 60 3.25 14.47 69 59 3.43
CONSOLE [34] 13.59 73 63 3.00 14.31 72 61 3.30
ESCENE [62] 10.80 68 64 3.39 11.89 66 63 3.77
DUET [6] 13.94 72 60 3.31 14.73 69 59 3.65
SUSA (Ours) 12.18 73.05 64.85 3.06 13.27 72.56 63.83 3.23
Table 2: Comparison with state of the art on the R2R dataset. Bold and Underline highlight the best and the runner-up performance in each column. $\downarrow$ indicates better performance with lower values.* expressing our reimplementation results.
SOON. On the SOON dataset, as shown in Tab. 3, SUSA achieves state-of-the-art performance across multiple metrics, including SR, SPL, and RGSPL, attaining SPL scores of 30.82% and 25.47% on the validation and test unseen splits, respectively. Moreover, SUSA consistently enhances the baseline (DUET)’s navigation performance and generalization on three datasets, underscoring the effectiveness of integrating textual semantics and depth exploration maps.
Method Validation Unseen Test Unseen
[…]
GridMM [56] 38.92 37.46 24.81 3.91 46.20 36.27 21.25 4.15
SEAT [53] - 36.87 24.87 3.91 - 35.89 22.55 4.47
KERM [31] 35.83 38.05 23.16 4.40 - - - -
DUET [6] 36.20 36.28 22.58 3.75 41.83 33.44 21.42 4.17
SUSA(Ours) 36.86 42.89 30.82 6.48 36.13 36.87 25.47 5.93
Table 3: Comparison with the state of the art on the SOON.
@@ 4.3 Diagnostic Experiment
[…]
ViT $w/o$ 51.07 46.98 33.73 32.15 23.03
CLIP $w/o$ 57.17 51.55 35.75 35.19 24.40
CLIP ResNet-50(ImageNet) 54.42 50.18 34.63 34.39 23.66
CLIP ResNet-50(Gibson) 56.72 52.00 39.21 35.05 26.54
Table 4: Ablation of different visual representations on the REVERIE validation unseen split. Gibson and ImageNet serve as pre-training sources for the depth encoder in the DSP module.
@@
Table 5: Ablation on the relative importance of static and dynamic matching in the TSU module, balanced by the balance factor $\delta$.
@@
1) Image Features vs. Structural Features. We analyze RGB and depth features contributions within the DSP module (Tab. 4). For RGB images, substituting ViT [11] features with those from CLIP [44] yields marked improvements across all metrics, highlighting that CLIP achieves a better understanding of the environment than its ViT-based baseline counterpart. As for depth features, we leveraged depth features extracted by ResNet-50 [17] pre-trained on the ImageNet [9] or Gibson [57] datasets to investigated the effect of depth exploration maps. Compared to ImageNet, the Gibson simulator provides spatial structural information that serves as valuable navigation priors. Therefore, in  Tab. 4, the ResNet-50 pre-trained on the Gibson extracts structural features that refine spatial perception, yielding notable improvements in the SPL metric without compromising SR or RGS. These improvements indicate that better spatial perception enhances navigation efficiency.
@@
2) Static Matching vs. Dynamic Matching. To assess the impact of static matching ($\delta=0$) and dynamic matching ($\delta=1$) strategies in the TSU module, we perform ablation studies based on balance factor $\delta$ on the R2R and REVERIE.  Tab. 5 summarizes these results, where w/o indicates the exclusion of the TSU module, corresponding to the last row results in Tab. 4. Furthermore, adaptive denotes that $\delta$ is a learnable parameter which adaptively adjusts the contribution of each matching strategy. Compared to the first row in Tab. 5, overall, incorporating the TSU module (regardless of static or dynamic matching strategy) improves SR and RGS, albeit at the cost of SPL on REVERIE. In contrast, it enhances SPL on R2R. While the TSU module can achieve superior performance under the adaptive setting, its performance is inferior when $\delta$ = 0.5. Specifically, at $\delta$ = 0.5, the SPL metric reaches 39.50% and 64.85% on the two datasets, with the RGS, nDTW, and sDTW metrics also achieving optimal performance. Excellent results demonstrate dual merits of proposed TSU module, which boosts object identification accuracy on the REVERIE task while simultaneously improving the fidelity between navigation trajectories and instructions on the R2R.
@@
Figure 5: a) and b) show the performance gap ($\downarrow$) between seen and unseen environments, while c) and d) present key metrics ($\uparrow$) for different pretraining strategies on the R2R and REVERIE.
3) Improved Generalization via Pretraining. As aforementioned in Sec. 3.5, pretraining the entire SUSA framework on multiple auxiliary tasks may lead to overfitting, which may negatively impact the generalization in unseen environments, particularly considering the modality differences between textual panorama, depth exploration maps, and RGB representations. As shown in Fig. 5, we observe that the partial pretraining strategy not only consistently narrow the performance gap between seen and unseen environments but also outperform complete/random pretraining strategies in navigation performance. For example, as depicted in Fig. 5 (a), the partial pretraining model achieves an SR gap of only 21.11% between the seen and unseen splits on REVERIE, reducing the gap by 5.92% compared to the complete pretraining strategy. Additionally, as shown in Fig. 5 (c), multiple metrics indicate that partial pretraining results in superior performance. Fig. 5(b) and (d) illustrate similar trends are observed on R2R, further emphasizing the generalization of the partial pretraining strategy.
ID DSP TSU HRA SR SPL RGSPL
1 46.98||||52.51 33.73||||36.06 23.03||||22.06
2 52.00||||52.53 39.21||||39.53 26.54||||24.97
3 53.17||||53.89 38.31||||40.89 26.73||||26.43
4 55.07||||51.97 39.50||||37.87 26.70||||25.20
5 51.75||||54.39 38.86||||41.54 26.56||||27.31
@@
Table 6: Ablation study on different components of SUSA on the REVERIE (validation||||test) unseen split.
@@
4) Overall Design. To thoroughly evaluate the efficacy of key components, we conduct ablation experiments on the REVERIE validation and test unseen split in the Tab. 6. As shown in Tab. 6, an intriguing phenomenon is observed: directly incorporating the DSP and TSU modules (#4) results in an improvement on validation unseen (e.g., SR=55.07, SPL=39.50), while performance on test unseen deteriorates. Conversely, contrastive learning within the HRA module (#5) produces opposite outcomes, particularly with respect to the SR metric. This may be attributed to the learnable token $v$, which, while pooling hybrid environmental representations, also introduces slight noise that causes fluctuations of navigation performance. Nevertheless, the overall performance is progressively improved with the inclusion of DSP, TSU and HRA modules. Analysis of the HRA module and the computational efficiency of the entire model is presented in supplementary materials.

### v5

@@ 3.4 Hierarchical Aggregation and Prediction
Hybrid Representations Aggregation. As depicted in Fig. 4, we employ attention pooling to project all environmental representations at step $t$, along with the corresponding instruction, into an equivalent dimension $d$. Specifically, the attention pooling AP operation is defined as follows:
$\hat{\mathcal{H}}=\text{AP}(\mathcal{H},v),\quad\mathcal{H}\in\{\tilde{\mathcal{X}^{\mathcal{S}}},\tilde{\mathcal{T}^{\mathcal{D}}},\tilde{\mathcal{X}^{R}},\tilde{\mathcal{T}^{R}}$$\mathcal{I}\},$ (4)
where $v\in\mathbb{R}^{d}$ denoting the learnable query vector and $\hat{\mathcal{H}}\in\mathbb{R}^{d}$. The AP operator is given by:
@@
[…]
$E^{\mathrm{hyb}}=\begin{bmatrix}\hat{\mathcal{X}^{\mathcal{S}}}&\hat{\mathcal{X}^{\mathcal{R}}}&\hat{\mathcal{T}^{\mathcal{D}}}&\hat{\mathcal{T}^{\mathcal{R}}}\end{bmatrix}\cdot\mathcal{B}$, (7)
where $\sigma$ stands for the sigmoid activation function, and $[;]$ denotes concatenation.
@@
We convert local predictions into global actions and predict $a_{t}$ via dynamic fusing (Chen et al. 2022b) $P_{\mathcal{T}}$ and $P_{\mathcal{X}}$.
3.5 Training
Partially Pretraining. Our pre-training phase followed prior work (Chen et al. 2022b), utilizing auxiliary tasks such as Masked Language Modeling (MLM), Masked Region Classification (MRC), and Single-step Action Prediction (SAP). Object Grounding (OG) task to localize target objects in REVERIE and SOON. Given the sparsity of depth images and textual semantics compared to RGB images, completely pre-training all parameter of SUSA may lead to overfitting and increased training costs. As shown in Fig. 2, only the computation flows indicated by black arrows were pre-trained. The effectiveness of this partial pre-trained strategy is demonstrated in Section 4.3.
Methods REVERIE R2R
Val Unseen Test Unseen Val Unseen Test Unseen
SR SPL RGSPL SR SPL RGSPL $\downarrow SR SPL NE\downarrow$
RCM (Wang et al. 2019) 9.2 6.9 3.8 7.8 6.6 3.1 43 - 6.0 43 38 6.1
GridMM (Wang et al. 2023c) 51.3 36.4 24.5 53.1 36.6 23.4 75 64 - 73 62 -
KERM (Li et al. 2023b) 49.0 34.8 24.1 52.2 37.4 23.1 71.9 60.9 3.2 69.7 59.2 3.6
AZHP (Zhan et al. 2024) 49.0 36.2 24.1 52.5 36.1 22.5 71 60 3.2 69 59 3.4
FDA (He et al. 2024) 47.5 35.9 24.3 49.6 36.4 22.0 72 64 3.4 69 62 3.41
CONSOLE (Lin et al. 2024) 50.0 34.4 23.3 55.1 37.1 22.2 73 63 3.0 72 61 3.3
ACME (Wu et al. 2025) 49.5 32.4 24.0 51.9 34.7 23.6 72.8 62.3 3.12 70.4 61.2 3.68
DUET (Chen et al. 2022b) 46.9 33.7 23.0 52.5 36.0 22.0 72 60 3.3 69 59 3.65
SUSA (Ours) 51.7 38.8 26.5 54.3 41.5 27.3 73.0 64.8 3.0 72.5 63.8 3.20
Table 1: Comparison with the state of the art on REVERIE and R2R. Bold highlights the best performance in each column. $\downarrow$ indicates better performance with lower values. “–”: unavailable statistics. % is omitted for brevity.
Contrastive and Imitation Learning. During fine-tuning, we employ a contrastive learning loss $\mathcal{L}_{cl}$ to align the hybrid environmental representations $E^{hyb}$ with the corresponding instructions $\mathcal{I}$:
@@ 3.5 Training
where $B$ and $t$ represent the batch size and the temperature. Then, we train the model based on hybrid environment representations by performing single-step action prediction (Chen et al. 2021, 2022b) with a contrastive learning loss $\mathcal{L}_{cl}$. The hybrid single action prediction loss in imitation learning is computed as $\mathcal{L}_{hsap}^{gt}=\sum_{t=1}^{T}-\log p(a_{t}^{gt}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$ and $\mathcal{L}_{hsap}^{\star}=\sum_{t=1}^{T}-\log p(a_{t}^{\star}|\mathcal{I},\mathcal{S}_{t},\mathcal{D}_{<t},\mathcal{R}_{<t})$, where $a_{t}^{gt}$ represents the ground truth action, and $a_{t}^{\star}$ denotes the pseudo-target action sampled from current exploration map. The final training loss is then expressed as follows:
@@
Method Val Unseen Test Unseen
SPL RGSPL SPL RGSPL
GridMM (Wang et al. 2023c) 24.8 3.9 21.2 4.1
KERM (Li et al. 2023b) 23.1 4.4 - -
SEAT (Wang et al. 2024b) 24.8 3.9 22.5 4.4
ACME (Wu et al. 2025) 27.8 4.3 22.3 5.5
DUET (Chen et al. 2022b) 22.5 3.7 21.4 4.1
SUSA(Ours) 30.8 6.4 25.4 5.9
Table 2: Comparison with the state of the art on the SOON.
Methods Val Unseen Test Unseen
SR SPL SR SPL
GridMM (Wang et al. 2023c) 49 41 46 39
AO-Planner (Chen et al. 2025) 47 33 - -
ETPNav (An et al. 2024) 57 49 55 48
DUET (Chen et al. 2022b) 51.9 41.2 47.1 40.5
SUSA (Ours) 52.7 43.6 50.9 43.9
Table 3: Comparison with the state of the art on R2R-CE. We directly transfer the DUET and SUSA as planners of ETPNav into the continuous environment without pretraining. Underline indicates improvements over the baseline.
@@ 4.1 Task Setup and Implementation
Datasets. We mainly evaluate our model on three diverse VLN benchmarks. R2R (Anderson et al. 2018) focuses solely on navigation following detailed instructions. REVERIE (Qi et al. 2020) requires the agent to recognize the correct object from candidate bounding boxes upon reaching the navigation goal. SOON (Zhu et al. 2021) challenges agents to generate candidate bounding boxes using an object detector. Given that discrete agents can serve as planners of ETPNav (An et al. 2024), we also evaluate SUSA on the continuous R2R-CE (Krantz et al. 2020). R2R-CE is constructed from the discrete Matterport3D and executed in the continuous Habitat simulator (Savva et al. 2019).
Metrics. We evaluate the navigation performance using standard metrics: Success Rate (SR), Oracle Success Rate (OSR), Navigation Error (NE), and Success Rate weighted by Path Length (SPL). For the REVERIE and SOON tasks, which involve object identification, we also evaluate object grounding metrics: Remote Grounding Success (RGS) and RGS weighted by Path Length (RGSPL). Given that SPL optimally balances navigation success rate and trajectory length, we employ it as our primary metric.
Implementation Details. Pre-training was performed for 100k iterations with a batch size of 32. Fine-tuning was performed for 25k iterations for VLN tasks. For fine-tuning, we used batch sizes of 4, 8, 2 and 16 for R2R, REVERIE, SOON and R2R-CE, respectively. The default hyperparameters are $\lambda_{1}$ = 0.2 and $\lambda_{2}$= 0.8. The node and view dimensions are identical: $d_{v}=d_{n}=d=768$. We minimized changes to the baseline (DUET) (Chen et al. 2022b) settings and used no additional annotations. All experiments were conducted on a single NVIDIA RTX 4090 GPU.
@@ 4.2 Comparisons with State of the Art
We primarily conduct a comprehensive comparison on three discrete VLN tasks—REVERIE (Qi et al. 2020), R2R (Anderson et al. 2018), and SOON (Zhu et al. 2021). For fairness, we exclude methods that involve pre-exploration (Sigurdsson et al. 2023) or utilize large-scale data augmentation (Wang et al. 2023b). Tables 1 and 2 present a comparative analysis on REVERIE, R2R, and SOON. Our method achieves, and in some cases surpasses, the performance of previous state-of-the-art approaches on various metrics. For instance, SUSA achieves an SPL/RGSPL of 41.5%/27.3% on the REVERIE test unseen split, significantly outperforming the baseline (DUET) by a large margin of 5.5%/5.3%.
Furthermore, we briefly transfer the agent into continuous environments to coarsely assess its adaptability and generalization. Results in Table 3 indicate that SUSA achieves superior performance over the baseline in the continuous R2R-CE task, highlighting the potential of hierarchical representation enrichment in continuous navigation scenarios.
@@ 4.3 Diagnostic Experiment
RGB Depth SR SPL RGS RGSPL
ViT $w/o$ 46.9 33.7 32.1 23.0
CLIP $w/o$ 51.5 35.7 35.1 24.4
CLIP ImageNet 50.1 34.6 34.3 23.6
CLIP Gibson 52.0 39.2 35.0 26.5
Table 4: Ablation of different visual representations on the REVERIE validation unseen split. Gibson and ImageNet serve as pre-training sources for the depth encoder (ResNet-50) in the DSP module.
@@
Table 5: Ablation of static and dynamic matching strategies in the TSU module, balanced by the balance factor $\delta$.
@@
1) Image Features vs. Structural Features. We analyze RGB and depth features contributions within the DSP module (Table  4). For RGB images, substituting ViT (Dosovitskiy et al. 2020) features with those from CLIP (Radford et al. 2021) yields marked improvements across all metrics. As for depth features, we leveraged depth features extracted by ResNet-50 (He et al. 2016) pre-trained on the ImageNet (Deng et al. 2009) or Gibson (Xia et al. 2018) datasets to investigated the effect of depth exploration maps. Compared to ImageNet, the Gibson simulator provides spatial structural information that serves as valuable navigation priors. Therefore, the ResNet-50 pre-trained on the Gibson yields notable improvements in the SPL metric without compromising SR or RGS. These improvements indicate that better spatial perception enhances navigation efficiency.
@@
2) Static Matching vs. Dynamic Matching. To assess the impact of static matching ($\delta=0$) and dynamic matching ($\delta=1$) strategies in the TSU module, we perform ablation studies based on balance factor $\delta$ on the R2R and REVERIE. Table 5 summarizes these results, where w/o indicates the exclusion of the TSU module. Furthermore, adaptive denotes that $\delta$ is a learnable parameter which adaptively adjusts the contribution of each matching strategy. While the TSU module can achieve superior performance under the adaptive setting, it underperforms when $\delta$ = 0.5. Specifically, at $\delta$ = 0.5, the SPL reaches 39.50% and 64.85% on the two datasets. Excellent results demonstrate dual merits of TSU module, which boosts object identification accuracy on the REVERIE while simultaneously improving the fidelity between navigation trajectories and instructions on the R2R.
@@
Figure 5: (a) and (b) show the performance gap ($\downarrow$, lower values indicate better performance.) between seen and unseen environments, while (c) and (d) present key metrics ($\uparrow$) for different pretraining strategies on the R2R and REVERIE.
ID DSP TSU HAP SR SPL RGSPL
1 46.9||52.5 33.7||36.0 23.0||22.0
2 52.0||52.5 39.2||39.5 26.5||24.9
3 53.1||53.8 38.3||40.8 26.7||26.4
4 55.0||51.9 39.5||37.8 26.7||25.2
5 51.7||54.3 38.8||41.5 26.5||27.3
@@
Table 6: Ablation study on different components of SUSA on the REVERIE (validation||test) unseen split.
3) Improved Generalization via Pretraining. As aforementioned in Section 3.5, pretraining the entire SUSA framework on multiple auxiliary tasks may lead to overfitting, which negatively impact the generalization in unseen environments. As shown in Fig. 5, the partial pretraining strategy not only consistently narrows the performance gap between seen and unseen environments but also outperforms complete/random pretraining strategies in navigation performance. In Fig. 5 (a), the partial pretraining model achieves an SR gap of only 21.11% between the seen and unseen splits on REVERIE, reducing the gap by 5.92% compared to the complete pretraining strategy. Additionally, as depicted in Fig. 5 (c), multiple metrics indicate that partial pretraining results in superior performance. Fig. 5 (b) and (d) illustrate similar trends are observed on R2R, further emphasizing the generalization of the partial pretraining strategy.
@@
4) Overall Design. To thoroughly evaluate the efficacy of key components, we conduct ablation experiments on the REVERIE validation and test unseen splits in Table 6. An intriguing phenomenon is observed in Table 6: directly incorporating the DSP and TSU modules (#4) results in an improvement on validation unseen (e.g., SR=55.07, SPL=39.50), while performance on test unseen deteriorates. Conversely, contrastive learning within the HAP module (#5) produces opposite outcomes, particularly with respect to the SR metric. This may be attributed to the learnable token $v$, which, while pooling hybrid environmental representations, also introduces slight noise that causes fluctuations of navigation performance. Nevertheless, the overall performance is progressively improved with the inclusion of DSP, TSU and HAP modules.